A $26,000 loan at 8% compounded quarterly is to be repaid by equal quarterly payments over a seven-year term.

a.

What will be the principal component of the sixth payment? (Round your answer to 2 decimal places.)

Principal component of the sixth payment $ .
b.

What will be the interest portion of the twenty-second payment? (Round your answer to 2 decimal places.)

Interest portion $ .
c.

How much will the loan’s balance be reduced by Payments 10 to 15 inclusive? (Round your answer to 2 decimal places.)

Principal paid by Payment 10 to Payment 15 inclusive $ .
d.

How much interest will be paid in the second year? (Round your answer to 2 decimal places.)

Interest paid in the second year $ .
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

SOLVED WITH BA II PLUS CALCULATOR

C W ENG 04:13 01-08-2020 66 34 Х F1831 . B A ח с E I D G J K 1831 1832 1833 1834 1835 1836 set C/Y = P/Y = 4 BA II PLUS CALCU

C w ENG 04:13 01-08-2020 66 34 Х H1876 . A B с D I E F. G J K LOAN BALANCE AFTER 21 PAYMENTS N = 7 [28-21] 1/Y = 8.00 PMT = -

C W ENG 04:13 01-08-2020 66 34 Х H1903 . A B с I D E F. G J K 1880 1881 1882 1883 1884 LOAN BALANCE AFTER 4 PAYMENTS N = 24 [